OpenStudy (anonymous):

Can someone explain completing the square? x^2+y^2+12x-14y-36=0

OpenStudy (anonymous):

i ordered the equation x^2+12x+y^2-14y=36

OpenStudy (anonymous):

dont know what to do from here

OpenStudy (anonymous):

First get the 36 to the right side of the equation. Now what do you have?

OpenStudy (anonymous):

Sorry see you did that>

OpenStudy (anonymous):

Now we need to set the equation up. To complete the square in the equation ax^2+bx+c we take 1/2 of b and square it and then add to both sides of the equation. Then do the same for ay^2+by+c. Do you understand?

OpenStudy (anonymous):

so half of 12?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay i got x^2+12x+36+y^2-14y-49=36+36-49

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

Very good.. Now we need to factor the left side and combine the left side.

OpenStudy (anonymous):

(x+6)^2+(y-7)^2=23

OpenStudy (anonymous):

You go girl....very good

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

oh, if you still see this. I got 23, but webworks says 23 isn't correct? i dont see whats wrong